# Fuzzy Ontology Embeddings and Visual Query Building for Ontology Exploration

### v2

@@ 6 Conclusion and Future Work
This paper introduced FuzzyVis, a prototype system for visual query building over ontologies using fuzzy logic and membership-based embeddings. FuzzyVis allows users to define complex, possibly vague concepts using familiar set operations such as intersection, union, and negation. These user-defined concepts are interpreted as fuzzy sets and embedded in a continuous vector space, enabling efficient similarity-based query evaluation. The system supports exploratory ontology search through an interactive visual interface that combines search, navigation, query construction, and result presentation. FuzzyVis demonstrates how fuzzy embeddings can enhance user interaction with large ontologies, making query formulation more accessible and intuitive, especially for non-expert users. Unlike traditional querying methods that rely on formal syntax or strict logical reasoning, FuzzyVis bridges symbolic and sub-symbolic techniques to support approximate reasoning and flexible concept retrieval.
FuzzyVis is a prototype system, and there are several directions for future development. First, improving system performance and scalability for very large ontologies remains a priority. This includes optimizing both front-end visualizations and back-end embedding computations. Second, expanding FuzzyVis’s features can enhance its usability. For example, incorporating additional visualizations such as collapsible tree views can support different user needs. Similarly, adding new views for triaging and organizing query results would improve the analysis and comparison of retrieved concepts. Third, exploring alternative embedding techniques—especially those that are more scalable or tailored to specific ontology types—may lead to better performance or interpretability. Finally, an important research direction is to extend FuzzyVis to support more expressive ontologies with richer axioms beyond hierarchical subsumption structures. This would require augmenting both the query builder and embedding mechanism to handle a broader range of logical constructs and relationships.

### v3

@@ 6 Conclusion and Future Work
This paper presented FuzzyVis, a prototype system for visual query building over ontologies using fuzzy logic and membership-based embeddings. FuzzyVis enables users to define complex, potentially vague concepts through familiar set operations—intersection, union, and negation—and interprets these as fuzzy sets embedded in a continuous vector space. This representation supports efficient, similarity-based query evaluation. Through an interactive visual interface, FuzzyVis integrates search, navigation, query construction, and result presentation, making query formulation more accessible and intuitive, particularly for non-expert users. By combining symbolic and sub-symbolic techniques, the system supports approximate reasoning and flexible concept retrieval, in contrast to traditional methods that rely solely on formal syntax and strict logic.
As a prototype, FuzzyVis offers several avenues for future development. First, improving performance and scalability for very large ontologies remains essential, requiring optimization of both front-end visualizations and back-end embedding computations. Second, expanding functionality—such as adding collapsible tree views or new panels for organizing and comparing query results—could better support diverse user needs. Third, investigating alternative embedding techniques, especially those optimized for specific ontology types or designed for greater interpretability, may yield improved results. Finally, extending FuzzyVis to handle more expressive ontologies with richer axioms beyond hierarchical subsumption would require enhancements to both the query builder and embedding mechanism, enabling support for a broader range of logical constructs and relationships.